OpenStudy (anonymous):

Graph each function using the given values of x. Use the graph to determine the domain and range of the function. 1a. f(x) = x2 + 3; x = -3, -2, -1, 0, 1, 2, 3 1b. g(x) = ; x = -2, -1, 2, 7 1c. h(x) = |x + 2|; x = 0, 1, 2, 3, 4, 5, 6 1d. f(x) = -5; x = -3, -2, -1, 0, 1, 2, 3 1e. g(x) = x3 + 4; x = -2, -1, 0, 1, 2 Ok so I am suppose to help someone solve these and I don't know how, i basically need like an answer key lol but I need to know by tmrw and am sorta clueless, I don't even remember learning this lol

OpenStudy (anonymous):

or can someone just show me how to solve these, please!!!!!!!!!!

OpenStudy (anonymous):

@jim_thompson5910 @Luigi0210 @Loser66 any of you guys remember these?

OpenStudy (luigi0210):

all of them? o.O

OpenStudy (anonymous):

just one and I will do the rest by ur example

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

cuz tmrw I need to sorta have an idea how to solve these

OpenStudy (anonymous):

here is the easiest one \[ f(x) = -5; x = -3, -2, -1, 0, 1, 2, 3\] \(f(-3)=-5, f(-2)=-5, f(-1)=-5, f(0)=-5, f(1)=-5,f(2)=-5, f(3)=-5\) points on the graph are \[(-3,-5),(-2,-5),(-1,-5),(0,-5),...\]

OpenStudy (anonymous):

\[1a. f(x) = x^2 + 3; x = -3, -2, -1, 0, 1, 2, 3 \] \[f(-3)=(-3)^2+3=9+3=12\] \[f(-2)=(-2)^2=4+3=7\] \[f(-1)=(-1)^2=1+3=4\] etc etc

OpenStudy (dan815):

just sub in values and see the value of the function and graph 

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

ty so much!!!!!!!!!!!!

OpenStudy (anonymous):

@luigi can we do 1b. and then we can compare answers?